OpenStudy (anonymous):

(PLEASE HELP ASAP) A bus went 318km farther than a car. The car went one third as far as the bus. How far did each vehicle travel? Choices for the cars distance, 147,151,159 (STEPS NEEDED)

OpenStudy (aum):

Let b be the distance the bus went and c be the distance the car went. "bus went 318km farther than a car" b = c + 318 ------ (1) "car went one third as far as the bus" c = b / 3 multiply both sides by 3 3c = b ------ (2) Solve for b and c using equations (1) and (2).

OpenStudy (anonymous):

How would you do the first step?

OpenStudy (aum):

Take b from (2) and put it in (1) and solve for c. Then solve for b.

OpenStudy (anonymous):

So 3c =c +318? right